What is the pH of a solution of 1.00L of water with 0.0973 moles of NaOH in it ?
Answer #1

Step1 : to calculate the molarity of NaOH

Molarity= 0.0973/1 (mol/L)

= 0.0973 M

Since NaOH is strong base dissociate completely

So [OH​​​​​-​​​​] = 0.0973 M

P​​​​​​OH = - log(0.0973) = 1.01

P​​​​​​H​​​ = 14 - 1.01 = 12.99